# The syzygy distinguisher

### v4

@@ Abstract
Since its introduction in 1978, this is the first time an analysis of the McEliece cryptosystem breaks the exponential barrier.
@@ 1 Introduction
Slightly more formally, a security proof for the system relies on two assumptions:
@@
Assumption 2, on the other hand, stands on a firm theoretical ground: the decoding problem for generic linear codes is known to be NP-hard [2]. As such, it is believed to resist the advent of quantum computers, which made the McEliece system a good candidate for postquantum cryptography. However, starting with Prange’s information set decoding algorithm [24], generic decoding methods saw continuous incremental improvements over time. Joint with technological progress in computational power, this eventually led to practical message recovery attacks against the McEliece system with its initially proposed parameter set. However, it appears that this weakness was only the result of a too optimistic choice of parameters. With the need for new standards for postquantum cryptography, an updated version named Classic McEliece was proposed, still relying on binary Goppa codes, but with more conservative parameter sets adapted to resist the best generic decoding attacks with some margin of safety. In the design rationale of this new system [4] one can find an impressive list of several dozens of papers on generic decoding algorithms, ranging over the last five decades. It is then observed that all these algorithms have complexity exponential in the error-correcting capability of the code. Better, the constant in this exponential is still the same as in Prange’s original result: all improvements remain confined in terms of lesser order! This gives a quite convincing feeling that we could possibly have reached the intrinsic complexity of cryptanalysis of this system.
Invalidating this belief, we present a new distinguisher for alternant and Goppa codes, i.e. a basic structural analysis of the McEliece cryptosystem, whose asymptotic complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms. Moreover, for given finite parameters, it is more efficient than previous distinguishers or key recovery methods, while not suffering from their strong regime limitations: in particular, it applies to the codes used in Classic McEliece.
@@
Likewise we claim that Theorem 2.8 of [9] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among $[7,4]$-codes, where we restrict to codes whose square is the whole space — otherwise we can use the square distinguisher. For such a code we always have $\dim(I_{2}({\mathsf{C}}))=\binom{4+1}{2}-7=3$. Let $Q_{1},Q_{2},Q_{3}$ be a basis of $I_{2}({\mathsf{C}})$. By definition, $Q_{1},Q_{2},Q_{3}$ do not satisfy linear relations with coefficients in $\mathbb{F}$, however they can satisfy relations whose coefficients are forms of degree $1$. Such relations are also called degree $3$ syzygies. They live in the kernel of the degree $3$ Macaulay matrix
@@
  • On the other hand, we estimate the Betti numbers of random codes in terms of raw code parameters (length, dimension, distance). This is done in section 4, partially relying on heuristics for which we do not have full proofs, but still providing experimental evidence, and sometimes also partial theoretical arguments.
@@ Acknowledgements
Restricting to homogeneous polynomials of degree $r$, we have a surjective map $S_{r}\longrightarrow{\mathsf{C}}^{\langle r\rangle}$, whose kernel we denote $I_{r}({\mathsf{C}})$. We then define the homogeneous coordinate ring of ${\mathsf{C}}$:
@@ 2 Minimal resolutions and graded Betti numbers
Let $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the $k$-dimensional polynomial ring over $\mathbb{F}$, graded by total degree. If $M_{0}$ is a finitely generated graded $S$-module, and $F_{0}$ is the free module on a minimal system of homogeneous generators of $M_{0}$, then the (first) syzygy module of $M_{0}$ is $M_{1}=\ker(\;F_{0}\longrightarrow M_{0}\;)$. Iterating this construction, we obtain a minimal resolution
@@
Thus the $F_{i}$ are of the form
@@
where $S(-j)$ is the free $S$-module of rank $1$ generated in degree $j$, so $S(-j)_{d}=S_{d-j}$; and $\beta_{i,j}$, the $(i,j)$-th graded Betti number of $M_{0}$, is the number of elements of degree $j$ in a minimal system of generators of $M_{i}$.
From now on let ${\mathsf{C}}$ be a $[n,k]$-code, and $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ its homogeneous coordinate ring. Then $M_{0}$ is a dimension $1$ Cohen-Macaulay quotient of $S$, and by the Auslander-Buchsbaum formula its minimal resolution has length $k-1$. So we have an exact sequence
@@
with $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$, $F_{0}=S$, $M_{1}=I({\mathsf{C}})$, hence $\beta_{0,0}({\mathsf{C}})=1$, $\beta_{0,j}({\mathsf{C}})=0$ for $j\neq 0$, and $\beta_{1,j}({\mathsf{C}})$ is the number of homogeneous polynomials of degree $j$ in a minimal system of generators of $I({\mathsf{C}})$.
$\begin{array}[]{c|cccccc}&0&1&\dots&i&\dots&k-1\\ \hline\cr\vdots&\vdots&\vdots&&\vdots&&\vdots\\ r&\beta_{0,r}&\beta_{1,r+1}&\dots&\beta_{i,r+i}&\dots&\beta_{k-1,r+k-1}\\ r+1&\beta_{0,r+1}&\beta_{1,r+2}&\dots&\beta_{i,r+i+1}&\dots&\beta_{k-1,r+k}\\ \vdots&\vdots&\vdots&&\vdots&&\vdots\\ \end{array}$
with null entries marked as “--” for readability. In our situation, as $F_{0}=S$, the 00-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
By the minimality property of the resolution, if for some $i$ and $D$ we have $\beta_{i,j}=0$ for all $j\leq D$, then also $\beta_{i+1,j}=0$ for all $j\leq D+1$. Applied recursively, this means all the upper-right quadrant of the Betti diagram defined by $\beta_{i,D}$ vanishes. In particular, as the evaluation map $S_{1}\longrightarrow C$ is an isomorphism, we see that $I({\mathsf{C}})$ is generated in degrees $\geq 2$, which implies:
@@
Hence the 00-th row of the Betti diagram is $(1,-,-,\dots)$.
@@
Figure 3: the $[23,12]_{2}$ Golay code ($d=7$, $d^{\perp}=8$)
@@
We prove (29) first. Let $\mathcal{G}$ be a minimal system of homogeneous generators of $M_{r-2}$, and for each $j$ let $B_{r-2,j}\subseteq M_{r-2,j}$ be the linear subspace generated by the degree $j$ elements of $\mathcal{G}$. Then, as $M_{r-2}$ is generated in degrees $\geq r-1$, we have $B_{r-2,r-1}=M_{r-2,r-1}$, and $B_{r-2,r}$ is a supplementary subspace to $S_{1}\cdot B_{r-2,r-1}=\operatorname{im}(\varphi_{r})$ in $M_{r-2,r}$. This proves (29).
@@
To make things even more explicit we introduce the following matrices. Assume ${\mathsf{C}}$ is given by a generator matrix $\mathbf{G}\in\mathbb{F}^{k\times n}$, with rows $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$. Choose monomial bases $\mathcal{M}_{1}=(X_{a})_{1\leq a\leq k}$, $\mathcal{M}_{2}=(X_{a}X_{b})_{1\leq a\leq b\leq k}$, and $\mathcal{M}_{3}=(X_{a}X_{b}X_{c})_{1\leq a\leq b\leq c\leq k}$ of $S_{1}$, $S_{2}$, $S_{3}$, ordered with respect to some monomial order.
@@
$\mathbf{G}_{2}\in\mathbb{F}^{\,\binom{k+1}{2}\,\times\,n}$ (33)
@@
Then $I_{2}({\mathsf{C}})$ is the left kernel of $\mathbf{G}_{2}$, and we choose a basis $\mathcal{B}_{2}$ of this space. Thus $\mathcal{B}_{2}$ consists of $\beta_{1,2}$ vectors, each of which has its entries indexed by $\mathcal{M}_{2}$.
@@
1:construct the matrix $\mathbf{G}_{2}$ according to Definition 1
2:compute a left kernel basis $\mathcal{B}_{2}$ of $\mathbf{G}_{2}$
@@
Sometimes we only control the syzygies of certain subideals of $I({\mathsf{C}})$, and from these we want to deduce information on the syzygies of the whole ideal. Intuitively, we expect syzygies between elements of the subideals could be seen as syzygies between elements of the ideal. Of course this fails in general, but it remains true if we restrict to the linear strand:
@@ 3 Lower bounds from the Eagon-Northcott complex
Under mild hypotheses, this complex is exact [9, Th. A.2.60 & Th. 6.4], so it defines a resolution of the quotient $R/\operatorname{im}(\bigwedge^{g}\mathbf{\Phi})$ defined by the ideal generated by the maximal minors of $\mathbf{\Phi}$.
@@
where we replace $Z^{(j)}_{r;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
@@
Let $f\geq g\geq d\geq 2$ be integers, and let $\mathbf{\Phi}\in S_{1}^{g\times f}$ be a matrix whose entries are homogeneous linear forms, i.e. elements of $S_{1}$. Recall from [9, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero. Let $V_{\mathbf{\Phi}}\subseteq S_{1}^{g}$ be the column span of $\mathbf{\Phi}$, so $\dim(V_{\mathbf{\Phi}})\leq f$, and set
@@
By Lemma 4 we can extend scalars to $\mathbb{F}_{q^{m}}$. By Lemma 5, $I_{2}({\mathsf{C}})$ then contains the $2\times 2$ minors of the $2\times f$ matrix $\mathbf{\Phi}$, and this matrix $\mathbf{\Phi}$ is $1$-generic by Lemma 6. By Proposition 4 applied $s$ times (with $g=d=2$), $I_{2}({\mathsf{C}}_{s})$ then contains the $2\times 2$ minors of a $2\times f_{s}$ matrix $\mathbf{\Phi}_{s}$, where $f_{s}\geq f-s$, and this matrix $\mathbf{\Phi}_{s}$ is $1$-generic also. Then by [9, §6B], the $2\times 2$ minors of $\mathbf{\Phi}_{s}$ are linearly independent, and its Eagon-Northcott complex is exact with only nonzero Betti numbers $\beta_{0,0}=1$ and $\beta_{r-1,r}=(r-1)\binom{f_{s}}{r}$ for $2\leq r\leq f_{s}$. We then conclude with Corollary 2 (with $l=1$ and $V^{(1)}=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi}_{s})$).
@@
It is possible to improve the lower bound on $\beta_{r-1,r}({\mathsf{C}}_{s})$, using not only $\mathbf{\Phi}=\mathbf{\Phi}^{(0)}$, but also its conjugates $\mathbf{\Phi}^{(1)},\dots,\mathbf{\Phi}^{(m-1)}$. See the corresponding discussion in the Supplementary material.

### v5

# & Télécom Paris, C2

@@ Abstract
Since its introduction in 1978, this is the first time an analysis111 we exclude results such as [28] that use an attack model for which direct countermeasures exist of the McEliece cryptosystem breaks the exponential barrier.
@@ 1 Introduction
Slightly more formally, a traditional security proof for the system relies on two assumptions:
@@
Assumption 2, on the other hand, stands on a firm theoretical ground: the decoding problem for generic linear codes is known to be NP-hard [2]. As such, it is believed to resist the advent of quantum computers, which made the McEliece system a good candidate for postquantum cryptography. However, starting with Prange’s information set decoding algorithm [25], generic decoding methods saw continuous incremental improvements over time. Joint with technological progress in computational power, this eventually led to practical message recovery attacks against the McEliece system with its initially proposed parameter set. However, it appears that this weakness was only the result of a too optimistic choice of parameters. With the need for new standards for postquantum cryptography, an updated version named Classic McEliece was proposed, still relying on binary Goppa codes, but with more conservative parameter sets adapted to resist the best generic decoding attacks with some margin of safety. In the design rationale of this new system [4] one can find an impressive list of several dozens of papers on generic decoding algorithms, ranging over the last five decades. It is then observed that all these algorithms have complexity exponential in the error-correcting capability of the code. Better, the constant in this exponential is still the same as in Prange’s original result: all improvements remain confined in terms of lesser order! This might give a feeling that we could possibly have reached the intrinsic complexity of cryptanalysis of this system.
Invalidating this belief, we present a new distinguisher for alternant and Goppa codes, i.e. a basic structural analysis of the McEliece cryptosystem, whose asymptotic complexity is subexponential in the error-correcting capability, hence better than that of generic decoding algorithms. Moreover, for given finite parameters, it is more efficient than state-of-the-art distinguishers or key recovery methods, while not suffering from their strong regime limitations: in particular, it applies to the codes used in Classic McEliece.
@@
Likewise we claim that Theorem 2.8 of [10] provides a $\beta_{2,3}$-based distinguisher for $\operatorname{{\mathsf{GRS}}}$ codes among $[7,4]$-codes,222erratum: among MDS $[7,4]$-codes (which then implies of full square) where we restrict to codes whose square is the whole space — otherwise we can use the square distinguisher. For such a code we always have $\dim(I_{2}({\mathsf{C}}))=\binom{4+1}{2}-7=3$. Let $Q_{1},Q_{2},Q_{3}$ be a basis of $I_{2}({\mathsf{C}})$. By definition, $Q_{1},Q_{2},Q_{3}$ do not satisfy linear relations with coefficients in $\mathbb{F}$, however they can satisfy relations whose coefficients are forms of degree $1$. Such relations are also called degree $3$ syzygies. They live in the kernel of the degree $3$ Macaulay matrix
@@
  • On the other hand, we estimate the Betti numbers of random codes in terms of raw code parameters (length, dimension, distance). This is done in section 4, partially relying on a natural heuristic: random codes are not expected to admit more syzygies than those forced by these parameters. We do not have full proofs for this fact, but we provide experimental evidence and partial theoretical arguments that support it.
@@
Restricting to homogeneous polynomials of degree $r$, we have a surjective map $S_{r}\longrightarrow{\mathsf{C}}^{\langle r\rangle}$, whose kernel we denote $I_{r}({\mathsf{C}})$. We then define the homogeneous coordinate ring of ${\mathsf{C}}$ as the formal direct sum
@@ 2 Minimal resolutions and graded Betti numbers
Let $S=\mathbb{F}[X_{1},\dots,X_{k}]$ be the $k$-dimensional polynomial ring over $\mathbb{F}$, graded by total degree. If $M_{0}$ is a finitely generated graded $S$-module, and $F_{0}$ is the free module on a minimal system of homogeneous generators of $M_{0}$, then the (first) syzygy module of $M_{0}$ is $M_{1}=\ker(\;F_{0}\longrightarrow M_{0}\;)$. Concretely, if $g_{1},\dots,g_{N}$ are minimal generators of $M_{0}$, elements of $F_{0}$ can be seen as formal sums $\sum_{u}f_{u}[g_{u}]$ with $f_{u}\in S$, where the $[g_{u}]$ are just formal symbols; such a formal sum then lies in $M_{1}$ when the actual sum evaluated in $M_{0}$ satisfies $\sum_{u}f_{u}g_{u}=0$.
Iterating this construction, we obtain a minimal resolution
@@
The number $\beta_{i,j}$ of degree $j$ elements in a minimal system of generators of $M_{i}$ does not depend on the choices made, and is called the $(i,j)$-th graded Betti number of $M_{0}$. Keeping track of the grading, we have
@@
where $S(-j)$ is the free rank $1$ module generated in degree $j$, so $S(-j)_{d}=S_{d-j}$.
$\begin{array}[]{c|ccccc}&0&1&\dots&i&\dots\\ \hline\cr\vdots&\vdots&\vdots&&\vdots&\\ r&\beta_{0,r}&\beta_{1,r+1}&\dots&\beta_{i,r+i}&\dots\\ r+1&\beta_{0,r+1}&\beta_{1,r+2}&\dots&\beta_{i,r+i+1}&\dots\\ \vdots&\vdots&\vdots&&\vdots&\\ \end{array}$
with null entries marked as “--” for readability.
Lemma 1
Assume $M_{i}$ is generated in degrees $\geq D$, i.e. $\beta_{i,j}=0$ for all $j\leq D-1$. Then $M_{i+1}$ is generated in degrees $\geq D+1$, and by induction all the upper-right quadrant of the Betti diagram defined by $\beta_{i,D-1}$ vanishes.
Proof
Let $g_{1},\dots,g_{N}$ form a minimal system of generators of $M_{i}$. By hypothesis all $g_{u}$ have degree $\geq D$. Then, in a homogeneous relation $\sum_{u}f_{u}g_{u}=0$, no $f_{u}$ can be a nonzero constant: otherwise the corresponding $g_{u}$ could be expressed as a linear combination of the others, hence could be removed from the system of generators, contradicting minimality. Thus all nonzero $f_{u}$ have degree $\geq 1$, and the element of $M_{i+1}$ defined by this homogeneous relation has degree $\geq D+1$.
As $M_{i}$ and $F_{i}$ are graded $S$-modules, we will also write $M_{i,j}$ and $F_{i,j}$ for their degree $j$ homogeneous components.
From now on let ${\mathsf{C}}$ be a $[n,k]$-code, with generator matrix $\mathbf{G}$, and $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ its homogeneous coordinate ring. Then $M_{0}$ is a dimension $1$ Cohen-Macaulay quotient of $S$, and by the Auslander-Buchsbaum formula its minimal resolution has length $k-1$. So we have an exact sequence
@@
and the Betti diagram has $k$ columns, indexed from 00 to $k-1$.
As $M_{0}={\mathsf{C}}^{\langle\cdot\rangle}$ and $F_{0}=S$, we have $\beta_{0,0}=1$ and $\beta_{0,j}=0$ for $j\neq 0$, i.e. the 00-th column of the Betti diagram is always $(1,-,-,\dots)^{\top}$.
Likewise $M_{1}=I({\mathsf{C}})$, and $\beta_{1,j}$ is the number of homogeneous polynomials of degree $j$ in a minimal system of generators of $I({\mathsf{C}})$. As the evaluation map $S_{1}\longrightarrow C$ is an isomorphism, we see that $I({\mathsf{C}})$ is generated in degrees $\geq 2$. Lemma 1 then implies:
@@
Hence the 00-th row of the Betti diagram is always $(1,-,-,\dots)$.
@@
Figure 3: the $[23,12]_{2}$ Golay code
Observe that replacing $\mathbf{G}$ with $\mathbf{S}\mathbf{G}$, for $\mathbf{S}\in\mathbb{F}^{k\times k}$ an invertible matrix, corresponds to a linear change of coordinates for the variables $X_{1},\dots,X_{k}$. This is an automorphism of $S$, hence will not affect the Betti numbers. Thus these Betti numbers really are invariants of ${\mathsf{C}}$, independently of the choice of $\mathbf{G}$.
More generally recall that two $[n,k]$ codes ${\mathsf{C}}_{1},{\mathsf{C}}_{2}$, defined by generator matrices $\mathbf{G}_{1},\mathbf{G}_{2}$, are linearly isometric, or monomially equivalent, if $\mathbf{G}_{1}=\mathbf{S}\mathbf{G}_{2}\mathbf{P}\mathbf{D}$ for $\mathbf{S}\in\mathbb{F}^{k\times k}$ an invertible matrix, and $\mathbf{P},\mathbf{D}\in\mathbb{F}^{n\times n}$ a permutation matrix and an invertible diagonal matrix.
Lemma 3
Monomially equivalent codes have the same Betti numbers.
Proof
We already discussed the action of $\mathbf{S}$. As for the action of $\mathbf{P}\mathbf{D}$, it preserves the kernel of the graded evaluation map $S\longrightarrow{\mathsf{C}}^{\langle\cdot\rangle}$.
@@
We prove (29) first. Let $\mathcal{G}$ be a minimal system of homogeneous generators of $M_{r-2}$, and for each $j$ let $B_{r-2,j}\subseteq M_{r-2,j}$ be the linear subspace generated by the degree $j$ elements of $\mathcal{G}$. Then, as $M_{r-2}$ is generated in degrees $\geq r-1$, we have $B_{r-2,r-1}=M_{r-2,r-1}$, and $B_{r-2,r}$ is a complementary subspace to $S_{1}\cdot B_{r-2,r-1}=\operatorname{im}(\varphi_{r})$ in $M_{r-2,r}$. This proves (29).
@@
To make things entirely explicit we now introduce the following matrices. Assume ${\mathsf{C}}$ is given by a generator matrix $\mathbf{G}\in\mathbb{F}^{k\times n}$, with rows $\mathbf{c}_{1},\dots,\mathbf{c}_{k}$. Choose monomial bases $\mathcal{M}_{1}=(X_{a})_{1\leq a\leq k}$, $\mathcal{M}_{2}=(X_{a}X_{b})_{1\leq a\leq b\leq k}$, and $\mathcal{M}_{3}=(X_{a}X_{b}X_{c})_{1\leq a\leq b\leq c\leq k}$ of $S_{1}$, $S_{2}$, $S_{3}$, ordered with respect to some monomial order.
@@
$\mathbf{M}_{2}\in\mathbb{F}^{\,\binom{k+1}{2}\,\times\,n}$ (33)
@@
Then $I_{2}({\mathsf{C}})$ is the left kernel of $\mathbf{M}_{2}$, and we choose a basis $\mathcal{B}_{2}$ of this space. Thus $\mathcal{B}_{2}$ consists of $\beta_{1,2}$ vectors, each of which has its entries indexed by $\mathcal{M}_{2}$.
@@
(Said otherwise, the rows of $\mathbf{M}_{3}$ are formed by multiplying all $\mathbf{q}\in\mathcal{B}_{2}$, seen as quadratic forms $\mathbf{q}=\sum_{1\leq b\leq c\leq k}q_{X_{b}X_{c}}X_{b}X_{c}$, by all variables $X_{a}$.)
@@
(Said otherwise, the rows of $\mathbf{M}_{r}$ are formed by multiplying all $\mathbf{s}\in\mathcal{B}_{r-1}$ by all variables $X_{a}$, where each $\mathbf{s}$ is seen as a formal sum $\mathbf{s}=\sum_{\mathbf{t}\in\mathcal{B}_{r-2}}L_{\mathbf{t}}[\mathbf{t}]$ whose coefficients are linear forms $L_{\mathbf{t}}=\sum_{1\leq b\leq k}s_{X_{b},\mathbf{t}}X_{b}$, resulting in formal sums $X_{a}\mathbf{s}=\sum_{\mathbf{t}\in\mathcal{B}_{r-2}}(X_{a}L_{\mathbf{t}})[\mathbf{t}]$ whose coefficients are quadratic forms.)
@@
1:construct the matrix $\mathbf{M}_{2}$ according to Definition 1
2:compute a left kernel basis $\mathcal{B}_{2}$ of $\mathbf{M}_{2}$
@@
Sometimes we only control the syzygies of certain subideals of $I({\mathsf{C}})$, and from these we want to deduce information on the syzygies of the whole ideal. Intuitively, we expect that syzygies between elements of the subideals could be seen as syzygies between elements of the ideal. Of course this fails in general, but it remains true if we restrict to the linear strand:
@@ 3 Lower bounds from the Eagon-Northcott complex
Under mild hypotheses, this complex is exact [10, Th. A.2.60 & Th. 6.4], so it defines a resolution of the quotient $R/\operatorname{im}(\bigwedge^{g}\mathbf{\Phi})$ defined by the ideal generated by the maximal minors of $\mathbf{\Phi}$.
@@
where notation $\widehat{i_{u}}$ means $i_{u}$ is omitted, and we replace $Z^{(j)}_{r;i_{1},\dots,i_{r}}$ with zero if $j\leq 0$ or $j\geq r$. Then, for all $r\geq 3$, we have
@@
Let $f\geq g\geq d\geq 2$ be integers, and let $\mathbf{\Phi}\in S_{1}^{g\times f}$ be a matrix whose entries are homogeneous linear forms, i.e. elements of $S_{1}$. Recall from [10, §6B] that such a matrix is said to be $1$-generic if for any nonzero $\mathbf{a}\in\overline{\mathbb{F}}^{g}$ and $\mathbf{b}\in\overline{\mathbb{F}}^{f}$, $\mathbf{a}\mathbf{\Phi}\mathbf{b}^{\top}$ is nonzero. Let $V_{\mathbf{\Phi}}\subseteq S_{1}^{g}$ be the column span of $\mathbf{\Phi}$, so $\dim(V_{\mathbf{\Phi}})\leq f$, and set
@@
By Lemma 6 we can extend scalars to $\mathbb{F}_{q^{m}}$. By Lemma 7, $I_{2}({\mathsf{C}})$ then contains the $2\times 2$ minors of the $2\times f$ matrix $\mathbf{\Phi}$, and this matrix $\mathbf{\Phi}$ is $1$-generic by Lemma 8. By Proposition 4 applied $s$ times (with $g=d=2$), $I_{2}({\mathsf{C}}_{s})$ then contains the $2\times 2$ minors of a $2\times f_{s}$ matrix $\mathbf{\Phi}_{s}$, where $f_{s}\geq f-s$, and this matrix $\mathbf{\Phi}_{s}$ is $1$-generic also. Then by [10, §6B], the $2\times 2$ minors of $\mathbf{\Phi}_{s}$ are linearly independent, and its Eagon-Northcott complex is exact with only nonzero Betti numbers $\beta_{0,0}=1$ and $\beta_{r-1,r}=(r-1)\binom{f_{s}}{r}$ for $2\leq r\leq f_{s}$. We then conclude with Corollary 2 (with $l=1$ and $V^{(1)}=\operatorname{im}(\bigwedge^{2}\mathbf{\Phi}_{s})$).
@@
It is possible to improve the lower bound on $\beta_{r-1,r}({\mathsf{C}}_{s})$, using not only $\mathbf{\Phi}=\mathbf{\Phi}^{(0)}$, but also its conjugates $\mathbf{\Phi}^{(1)},\dots,\mathbf{\Phi}^{(m-1)}$. See the corresponding discussion at the beginning of the Supplementary material.